### v1

@@ Abstract
In sensitive application domains, multi-objective hyperparameter selection can ensure the reliability of AI models prior to deployment, while optimizing auxiliary performance metrics. The state-of-the-art Pareto Testing (PT) method guarantees statistical reliability constraints by adopting a multiple hypothesis testing framework. In PT, hyperparameters are validated one at a time, following a data-driven order determined by expected reliability levels. This paper introduces a novel framework for multi-objective hyperparameter selection that captures the interdependencies among the reliability levels of different hyperparameter configurations using a directed acyclic graph (DAG), which is termed the reliability graph (RG). The RG is constructed based on prior information and data by using the Bradley-Terry model. The proposed approach, RG-based PT (RG-PT), leverages the RG to enable the efficient, parallel testing of multiple hyperparameters at the same reliability level. By integrating False Discovery Rate (FDR) control, RG-PT ensures robust statistical reliability guarantees and is shown via experiments across diverse domains to consistently yield superior solutions for multi-objective calibration problems.
@@ 1.1 Context and Motivation
Hyperparameter optimization is an essential element of the machine learning pipeline, particularly in sensitive domains such as healthcare, finance, and engineering Bischl et al. (2023); Yang & Shami (2020). Hyperparameters to be set prior to deployment include training parameters such as learning rates, regularization constants, and reward priorities in reinforcement learning systems Eimer et al. (2023), inference parameters, such as decision thresholds Hancock et al. (2022), and complexity settings determining numerical accuracy and computational efficiency Hu et al. (2021). Hyperparameter selection typically targets multiple performance objectives, such as classification accuracy, calibration, fairness, and domain-specific criteria, such as tracking precision Dong et al. (2018), perplexity Rahnama et al. (2018), and transmission rates Chen et al. (2020).
Figure 1: Comparison of Pareto Testing (PT) Laufer-Goldshtein et al. (2023) and reliability graph-based PT (RG-PT) proposed in this work. (a) PT first estimates the Pareto front in the space of risk functions to be controlled (horizontal axis) and optimized (vertical axis), with the red dashed line indicating the reliability constraint. (b) PT then tests the reliability of hyperparameters $\lambda$ sequentially, in the order of estimated reliability. (c) RG-PT first estimates interdependencies among the reliability levels of different hyperparameter configurations using a directed acyclic graph (DAG), which is termed the reliability graph (RG). This enables efficient parallel testing of multiple hyperparameters, potentially discovering more effective hyperparameter configurations (see Fig. 2).
Validation can be costly. Notably, determining training or fine-tuning parameters in large language models (LLMs) requires running expensive optimization procedures, and many engineering problems require interaction with the real world Polese et al. (2023). Thus, in modern settings, it is common to limit the search to a pre-selected discrete set of configurations for the hyperparameters Bergstra & Bengio (2012).
Most existing hyperparameter selection strategies make a best-effort attempt at addressing multi-objective problems. This is done by estimating functions of the Pareto front for the performance criteria of interest Deb et al. (2002); Hernández-Lobato et al. (2016); Paria et al. (2020); Li et al. (2020). An important, recent, exception is Pareto Testing (PT) Laufer-Goldshtein et al. (2023). PT builds on Learn-Then-Test (LTT) Angelopoulos et al. (2021), which pioneered the use of multiple hypothesis testing (MHT) to ensure the statistical validity of hyperparameter selection.
As illustrated in Fig. 1, PT carries out a preliminary best-effort optimization step to estimate the Pareto front in the space of the objectives of interest. These encompass risk functions that must be controlled, encoding reliability constraint, as well as auxiliary risk functions to be optimized. PT proceeds by testing the candidate hyperparameters according to the linear order determined by the increasing value of the estimated reliability risk functions (see Fig. 1(b)).
This paper is motivated by the observation that, especially in engineering applications, the hyperparameter space is often highly structured, and not all configurations can be neatly ordered by their reliability levels. To account for this, we introduce a novel multi-objective hyperparameter selection method that operates on a directed acyclic graph (DAG), termed the reliability graph (RG). In an RG, nodes represent candidate hyperparameter configurations and an edge directed from one hyperparameter to another indicates that the reliability of the first hyperparameter is predictive of the reliability level of the second hyperparameter (see Fig. 1(c)).
The proposed hyperparameter selection method, referred to as RG-based Pareto Testing (RG-PT), constructs an RG based on prior knowledge and limited validation data. Prior knowledge may include properties such as the reduction in reliability that may result from an increased sparsity at the levels of model parameters Zhou et al. (2021) or activations Chen et al. (2023). By incorporating MHT mechanisms operating on a DAG Ramdas et al. (2019) within the PT framework, RG-PT allows for a more efficient testing that can discover more effective hyperparameter configurations, while offering reliability guarantees (see Fig. 2).
@@ 1.2 Further Related Work
Hyperparameter Selection: Traditional methods for hyperparameter optimization like grid search and random search Bergstra & Bengio (2012) are widely used but they can be inefficient. In contrast, more advanced techniques such as Bayesian optimization Snoek et al. (2012) and bandit-based methods, such as Hyperband Li et al. (2018) and gradient-based optimization Maclaurin et al. (2015), can enhance computational efficiency via resource allocation, but they lack statistical guarantees. This gap addressed is by the mentioned frameworks incorporating MHT Angelopoulos et al. (2021); Laufer-Goldshtein et al. (2023).
Multi-Objective Optimization: Modern AI applications often require optimizing multiple objectives such as accuracy and efficiency. This can be formally done through Pareto optimization to identify all the feasible trade-off points among different objectives Deb et al. (2002). PT Laufer-Goldshtein et al. (2023) applies MHT to provide statistical reliability while balancing multiple objectives.
Conformal Prediction and LTT: Conformal prediction constructs statistically valid prediction intervals using calibration data Shafer & Vovk (2008); Romano et al. (2019). Drawing inspiration from conformal prediction, LTT Angelopoulos et al. (2021) frames hyperparameter selection as an MHT problem. While Angelopoulos et al. (2021) mentions the possible use of graph-based approaches, these are limited to fixed user-defined graphs or linear directed graphs (chains). Other extensions of LTT include quantile control Farzaneh et al. (2024), applications to information theory Farzaneh & Simeone (2024), and adaptive testing mechanisms Zecchin & Simeone (2024).
@@ 1.3 Main Contributions
Figure 2: Illustration of the benefits of the proposed RG-PT hyperparameter selection scheme over the state-of-the-art LTT and PT for an object detection application Angelopoulos et al. (2021). The red arrows mark the objects not detected by an object recognition model calibrated using LTT or PT that are instead detected by the same model calibrated via RG-PT (see Sec. 5.2 for details).
@@
Methodology: We propose RG-PT, a novel multi-objective hyperparameter selection framework that systematically utilizes interdependencies among the reliability levels of candidate hyperparameter configurations. RG-PT infers precedence relationships among the reliability levels of hyperparameters from prior information and data by constructing an RG via the Bradley-Terry (BT) ranking model Hunter (2004). Then, it applies MHT-based hyperparameter selection on the RG by following the graphical testing method in Ramdas et al. (2019).
Applications: We demonstrate the effectiveness of RG-PT through experiments in various scenarios, including language model calibration Peters & Martins (2021), object detection Angelopoulos et al. (2021), image classification Franceschi et al. (2024), and telecommunications engineering Valcarce (2020), highlighting its advantages over conventional methods.
The rest of this paper is organized as follows. In Sec. 2, we define the multi-objective hyperparameter selection problem. Sec. 3 provides an overview of PT, and Sec. 4 details the proposed RG-PT framework. Experimental results are presented in Sec. 5. We conclude the paper in Sec. 6.
@@ 2 Multi-Objective Hyperparameter Selection
In this section, we define the problem of multi-objective hyperparameter selection, and we show how this problem can be formulated via MHT by following references Angelopoulos et al. (2021); Laufer-Goldshtein et al. (2023).
@@ 2.1 Problem Definition
Consider a predefined discrete and finite set $\Lambda$ of hyperparameters $\lambda$, which govern the performance of a machine learning model (see, e.g., Bergstra & Bengio (2012)). Hyperparameter selection aims at obtaining a hyperparameter $\lambda\in\Lambda$ by considering its performance via $L$ risk functions. Specifically, when tested on a data point $Z$, a hyperparameter $\lambda$ attains risk values $r_{l}(Z,\lambda)$ for $l=1,\ldots,L$. The risk functions $r_{l}(Z,\lambda)$ are negatively oriented, meaning that lower risk values correspond to better-performing hyperparameters. The risks are normalized within the range $0\leq r_{l}(Z,\lambda)\leq 1$.
For each performance criterion $l=1,\ldots,L$, the average risk function is defined as
@@
2. Auxiliary risk functions: The second set of performance measures $\{R_{l}(\lambda)\}_{l=L_{c}+1}^{L}$ are unconstrained, and they are ideally optimized via the selection of the hyperparameter $\lambda$.
@@
which targets the minimization of the auxiliary risk functions $\{R_{l}(\lambda)\}_{l=L_{c}+1}^{L}$ under constraints on the reliability risk functions $\{R_{l}(\lambda)\}_{l=1}^{L_{c}}$. Solving a multi-objective optimization problem such as (3) may entail identifying the entire Pareto front of dominant solutions $\lambda\in\Lambda$, or obtaining specific solutions corresponding to scalar criteria Deb et al. (2016); Jamieson & Talwalkar (2016).
The problem (3) cannot be directly addressed since the data distribution $P_{Z}$ is unknown. However, we assume to have access to i.i.d. data $\mathcal{Z}=\{Z_{j}\}_{j=1}^{n}$ drawn from the unknown data distribution $P_{Z}$. For any subset $\widetilde{\mathcal{Z}}\subseteq\mathcal{Z}$, we write as
$\hat{R}_{l}(\lambda|\widetilde{\mathcal{Z}})=\frac{1}{|\widetilde{\mathcal{Z}}|}\sum_{Z\in\widetilde{\mathcal{Z}}}r_{l}(Z,\lambda)$ (4)
the corresponding empirical estimate of risk function $R_{l}(\lambda)$.
@@ 2.2 Hyperparameter Selection as Multiple Hypothesis Testing
As proposed in Angelopoulos et al. (2021), hyperparameter selection can be formally addressed as an MHT problem. Accordingly, for each hyperparameter $\lambda\in\Lambda$, we define the null hypothesis $\mathcal{H}_{\lambda}$ that hyperparameter $\lambda$ is not reliable, violating the constraints (2), i.e.,
@@
Thus, rejecting the null hypothesis $\mathcal{H}_{\lambda}$ implies that hyperparameter $\lambda$ is reliable, meeting all the constraints in (2). A rejection is also referred to as a discovery. A discovery is false if the selected hyperparameter $\lambda$ is actually unreliable, satisfying the null hypothesis $\mathcal{H}_{\lambda}$.
As discussed in the next section, by formulating hyperparameter selection as an MHT problem, we can leverage statistical tools that guarantee false discovery rate (FDR) requirements Casella & Berger (2024), ensuring that any selected configuration satisfies the risk constraints in (3) with a prescribed probability.
3 Background: Pareto Testing
In this section, we review PT Laufer-Goldshtein et al. (2023), a state-of-the-art MHT-based multi-objective hyperparameter selection method.
3.1 Statistical Reliability Constraints
Problem (3) cannot be directly solved since we only assume access to the empirical estimates. Following LTT Angelopoulos et al. (2021), using the data set $\mathcal{Z}$, PT finds a subset $\hat{\Lambda}_{\mathcal{Z}}\subseteq\Lambda$ of hyperparameters that satisfies the statistical constraint
$\mathrm{Pr}_{\mathcal{Z}}\big{[}R_{l}(\lambda)\leq\alpha_{l}$ for all $\text{for all }l=1,\ldots,L_{c},$ (6)
$\text{and for all }\lambda\in\hat{\Lambda}_{\mathcal{Z}}\big{]}\geq 1-\delta,$
where $0\leq\delta\leq 1$ is a user-specified probability of failure, and the probability is evaluated over the calibration data $\mathcal{Z}$. The condition (6), known as family-wise error rate (FWER) control, ensures that all the hyperparameters $\lambda$ in the selected set $\hat{\Lambda}_{\mathcal{Z}}$ are reliable with probability at least $1-\delta$.
3.2 Pareto Testing
PT Laufer-Goldshtein et al. (2023) leverages held-out data to select the testing order for addressing problems of the form (3). The procedure is summarized in Fig. 3, and it consists of the following steps:
Estimating the Pareto front for all risk measures: PT divides the calibration dataset $\mathcal{Z}$ into two disjoint subsets, $\mathcal{Z}_{\text{OPT}}$ and $\mathcal{Z}_{\text{MHT}}$. Using the data set $\mathcal{Z}_{\text{OPT}}$, it identifies the subset $\Lambda_{\text{OPT}}\subseteq\Lambda$ of hyperparameters that are on the Pareto front of the space of estimated risk measures $\{\hat{R}_{l}(\lambda|\mathcal{Z}_{\text{OPT}})\}_{l=1}^{L}$. This is done by addressing the multi-objective optimization problem (3) with the estimates $\{\hat{R}_{l}(\lambda|\mathcal{Z}_{\text{OPT}})\}_{l=1}^{L}$ in lieu of the true risks $\{R_{l}(\lambda|\mathcal{Z}_{\text{OPT}})\}_{l=1}^{L}$ by using any existing algorithm Laufer-Goldshtein et al. (2023).
Ordering the hyperparameters: PT creates a global linear ordering on the hyperparameters in the set $\Lambda_{\text{OPT}}$ by using the estimates $\{\hat{R}_{l}(\lambda|\mathcal{Z}_{\text{OPT}})\}_{l=1}^{L_{c}}$ of the reliability risk functions.
FWER-controlling MHT: Using the data set $\mathcal{Z}_{\text{MHT}}$, FWER-controlling MHT is carried out via fixed sequence testing (FST) Maurer (1995) by following the order found in the previous step. This results in the subset $\hat{\Lambda}_{\mathcal{Z}}$ of reliable hyperparameters that are guaranteed to satisfy the condition (6).
Addressing the multi-objective optimization problem: Given the subset $\hat{\Lambda}_{\mathcal{Z}}$, PT addresses the problem
To elaborate on steps ②-③, for each null hypothesis $\mathcal{H}_{\lambda}$ in (5), given a data set $\widetilde{\mathcal{Z}}\subseteq\mathcal{Z}$, define the $p$-value Rice (2006)
$p_{\lambda,l}(\widetilde{\mathcal{Z}})=e^{-2|\widetilde{\mathcal{Z}}|(\alpha_{l}-\hat{R}_{l}(\lambda|\widetilde{\mathcal{Z}}))_{+}^{2}}$ (8)
for $l=1,\ldots,L_{c}$, and the combined $p$-value
@@ 3.2 Pareto Testing
In step ②, hyperparameters in set $\Lambda_{\text{OPT}}$ are ordered from low to high based on the combined $p$-value $p_{\lambda}(\mathcal{Z}_{\text{OPT}})$, reflecting an ordering from most to least reliable hyperparameters according to data $\mathcal{Z}_{\text{OPT}}$. FST, a sequential MHT algorithm (see Appendix A), is then applied by using the combined $p$-values $p_{\lambda}(\mathcal{Z}_{\text{MHT}})$.
Figure 3: Illustration of the operation of PT Laufer-Goldshtein et al. (2023) for two risk functions $R_{1}(\lambda)$ and $R_{2}(\lambda)$. The objective is to control the risk $R_{1}(\lambda)$ to be below the threshold $\alpha$, while minimizing the risk $R_{2}(\lambda)$: ① Approximate the Pareto frontier using the estimates $\hat{R}_{1}(\lambda|\mathcal{Z}_{\text{OPT}})$ and $\hat{R}_{2}(\lambda|\mathcal{Z}_{\text{OPT}})$ to form the subset of hyperparameters $\Lambda_{\text{OPT}}$; ②-④ Apply a sequential FDR-controlling algorithm on the hyperparameters in $\Lambda_{\text{OPT}}$ to obtain the subset $\hat{\Lambda}_{\mathcal{Z}}$, and select the hyperparameter $\lambda^{*}$ that minimizes $\hat{R}_{2}(\lambda|\mathcal{Z}_{\text{OPT}})$.

### v2

@@ Abstract
The selection of hyperparameters, such as prompt templates in large language models (LLMs), must often strike a balance between reliability and cost. In many cases, structural relationships between the expected reliability levels of the hyperparameters can be inferred from prior information and held-out data – e.g., longer prompt templates may be more detailed and thus more reliable. However, existing hyperparameter selection methods either do not provide formal reliability guarantees or are unable to incorporate structured knowledge in the hyperparameter space. This paper introduces reliability graph-based Pareto testing (RG-PT), a novel multi-objective hyperparameter selection framework that maintains formal reliability guarantees in terms of false discovery rate (FDR), while accounting for known relationships among hyperparameters via a directed acyclic graph. Edges in the graph reflect expected reliability and cost trade-offs among hyperparameters, which are inferred via the Bradley-Terry (BT) ranking model from prior information and held-out data. Experimental evaluations demonstrate that RG-PT significantly outperforms existing methods such as learn-then-test (LTT) and Pareto testing (PT) through a more efficient exploration of the hyperparameter space.
@@ 1.1 Context and Motivation
Consider the problem of selecting prompt templates for a large language model (LLM)-based sentiment analysis task [1]. In this setting, the LLM receives a natural language prompt template along with a movie review as input, and the goal is to determine whether the sentiment expressed in the review is positive or negative. As illustrated in Fig. 1, the prompt templates $\lambda$ are chosen from a set of pre-determined choices $\Lambda$, and the objective is to identify prompt templates $\lambda$ that elicit consistently accurate responses across inputs [1].
Longer prompts often yield more reliable outputs [2]. However, they are also more costly to the end user when pay-per-token billing schemes are applied. This is commonly the case for enterprise software incorporating AI-driven analytics or hosted LLM endpoints via the LLM application programming interface (API) [3]. As exemplified in Fig. 1(a), this suggests that the reliability of different prompt templates follows a directed acyclic graph (DAG) structure with nodes closer to the roots corresponding to more costly prompt templates with a higher expected reliability.
Figure 1: Illustrative example for prompt engineering in LLM-based sentiment analysis: (a) Prompt template candidates in set $\Lambda$ have expected reliability levels that can be arranged on a reliability graph (RG), so that each parent prompt template is expected to be more reliable than its child prompts; (b) Distribution of the length of the shortest prompt templates identified by LTT [4], PT [5], and the proposed RG-PT for the Stanford Sentiment Treebank dataset [6] (see Sec. 4.1 for details).
An ideal prompt engineering scheme would apply hyperparameter selection methods capable of selecting prompt templates that are as short as possible while enduring formal reliability guarantees. Existing hyperparameter selection methods, however, either do not meet formal reliability requirements [7, 8], or cannot incorporate the structured knowledge encoded in a graph like the DAG in Fig. 1(a).
In particular, Learn-Then-Test (LTT) [4] pioneered the use of multiple hypothesis testing (MHT) for hyperparameter selection, providing formal guarantees on the reliability of the returned subset of hyperparameters. However, LTT cannot incorporate structured information about relative expected reliability levels of the hyperparameters. Pareto Testing (PT) [5] builds on LTT to address multi-objective optimization problems. Specifically, PT infers a global, linear ordering over hyperparameters from held-out data based on their expected relative reliability. Thus, PT cannot account for more complex structured relationships between candidate hyperparameters as in the example of Fig. 1(a).
This paper proposes a novel framework, reliability graph-based PT (RG-PT), that systematically captures and exploits interdependencies between hyperparameter configurations for hyperparameter selection. RG-PT models the hyperparameter space as a DAG, termed the reliability graph (RG). In an RG, nodes correspond to candidate hyperparameter configurations, such as prompt templates, and edges encode reliability relationships. If there is an edge from a hyperparameter $\lambda_{i}$ to another $\lambda_{j}$ in the RG, then $\lambda_{i}$ is expected to be more reliable than $\lambda_{j}$.
For the running example of prompt design, as shown in Fig. 1(b) (detailed in Sec. 4), RG-PT is seen experimentally to select shorter prompt templates than LTT and PT, while still satisfying formal guarantees in terms of false discovery rate (FDR). The FDR measures the fraction of unreliable prompt templates returned by the hyperparameter selection scheme. This advantage of RG-PT stems from its ability to encode rich structural relationships among prompt templates, so as to explore the hyperparameter space more efficiently during the MHT procedure.
@@ 1.2 Further Related Work
Hyperparameter Selection: State-of-the-art techniques for hyperparameter optimization, such as Bayesian optimization [9], bandit-based methods [8], and gradient-based optimization [10], provide satisfactory empirical performance, but they lack statistical guarantees. LTT addresses this gap by incorporating MHT in the hyperparameter selection process [4]. Extensions of LTT are surveyed in [11]. Note that, while reference [4] mentions the possible use of graph-based approaches, these are limited to fixed user-defined graphs or linear directed graphs (chains).
Multi-Objective Optimization: Modern AI applications often require optimizing multiple objectives such as accuracy, efficiency, and cost. This can be formally done through Pareto optimization to identify all the feasible trade-off points among different objectives [12]. PT [5] extends LTT to settings with multiple objectives, inferring a linear testing order in the hyperparameter space based on held-out data.
@@ 1.3 Main Contributions
Methodology: We propose RG-PT, a novel multi-objective hyperparameter selection framework that systematically infers and utilizes interdependencies among the expected reliability levels of candidate hyperparameter configurations. RG-PT first constructs a DAG, known as RG, based on prior information and held-out data via the Bradley-Terry (BT) ranking model [13] and the non-negative Lasso [14]. Then, it applies MHT-based hyperparameter selection on the RG by following DAGGER, a graphical testing method introduced in [15].
Applications: We demonstrate the effectiveness of RG-PT through experiments in LLM prompt engineering, sequence-to-sequence translation, object detection, image classification, and telecommunications, highlighting its advantages over existing methods.
The rest of this paper is organized as follows. In Sec. 2, we define the multi-objective hyperparameter selection problem. Sec. 3 details the proposed RG-PT framework. Experimental results are presented in Sec. 4. We conclude the paper in Sec. 5.
@@ 2 Multi-Objective Hyperparameter Selection
In this section, we define the problem of multi-objective hyperparameter selection, and we show how this problem can be formulated via MHT by following references [4, 5].
@@ 2.1 Problem Definition
Consider a predefined discrete and finite set $\Lambda$ of hyperparameters $\lambda$, which govern the performance of a machine learning model such as an LLM (see, e.g., [16]). The discrete set $\Lambda$ is populated in a preliminary pre-selection step [17, 18, 19] using methods including LLM judges [20] and continuous optimizers like Bayesian optimization [7] and Hyperband [8].
In a multi-objective setting with $L$ risk functions, when tested on a data point $Z$, a hyperparameter $\lambda$ attains risk values $r_{l}(Z,\lambda)$ for $l=1,\ldots,L$. The risk functions $r_{l}(Z,\lambda)$ are negatively oriented, meaning that lower risk values correspond to better-performing hyperparameters. The risks are normalized within the range $0\leq r_{l}(Z,\lambda)\leq 1$. Furthermore, for each performance criterion $l=1,\ldots,L$, the average risk function is defined as
@@
2. Auxiliary risk functions: The second set of performance measures $\{R_{l}(\lambda)\}_{l=L_{c}+1}^{L}$ are unconstrained, and are optimized in a best-effort fashion via the selection of the hyperparameter $\lambda$.
@@
which targets the minimization of the auxiliary risk functions $\{R_{l}(\lambda)\}_{l=L_{c}+1}^{L}$ under constraints on the reliability risk functions $\{R_{l}(\lambda)\}_{l=1}^{L_{c}}$. For example, in the setting of Fig. 1, we wish to minimize the prompt length, while ensuring a constraint on the accuracy of the LLM’s outputs.
Solving a multi-objective optimization problem such as (3) ideally entails identifying the entire Pareto front of dominant solutions $\lambda\in\Lambda$, or at least obtaining specific solutions corresponding to scalar criteria [21, 22]. However, the problem (3) cannot be directly addressed since the data distribution $P_{Z}$ is assumed to be unknown. Instead, we assume to have access to i.i.d. data $\mathcal{Z}=\{Z_{j}\}_{j=1}^{n}$ drawn from the unknown data distribution $P_{Z}$. For any data subset $\widetilde{\mathcal{Z}}\subseteq\mathcal{Z}$, the empirical estimate of risk function $R_{l}(\lambda)$ can be obtained as
$\hat{R}_{l}(\lambda|\widetilde{\mathcal{Z}})=\frac{1}{|\widetilde{\mathcal{Z}}|}\sum_{Z\in\widetilde{\mathcal{Z}}}r_{l}(Z,\lambda).$ (4)
@@ 2.2 Hyperparameter Selection as Multiple Hypothesis Testing
As proposed in [4], hyperparameter selection can be formally addressed as an MHT problem. Accordingly, for each hyperparameter $\lambda\in\Lambda$, we define the null hypothesis $\mathcal{H}_{\lambda}$ that hyperparameter $\lambda$ violates the reliability constraints (2), i.e.,
@@
Thus, rejecting the null hypothesis $\mathcal{H}_{\lambda}$ implies that hyperparameter $\lambda$ meets all the constraints in (2). A rejection is also referred to as a discovery. A discovery is false if the selected hyperparameter $\lambda$ is actually unreliable, satisfying the null hypothesis $\mathcal{H}_{\lambda}$.
Given a dataset $\widetilde{\mathcal{Z}}\subseteq\mathcal{Z}$, evidence against the reliability of each candidate hyperparameter $\lambda\in\Lambda$ can be measured by the p-value [23]
$p_{\lambda,l}(\widetilde{\mathcal{Z}})=\exp(-2|\widetilde{\mathcal{Z}}|(\alpha_{l}-\hat{R}_{l}(\lambda|\widetilde{\mathcal{Z}}))_{+}^{2})$ (6)
for each reliability risk function $l=1,\ldots,L_{c}$, yielding the combined p-value
@@
The quantity $p_{\lambda}(\widetilde{\mathcal{Z}})$ can be shown to be a valid p-value for the null hypothesis $\mathcal{H}_{\lambda}$ in (5). Therefore, thresholding the statistic (7) yields a reliability test that meets type-I error probability constraints [5].
Furthermore, by formulating hyperparameter selection as an MHT problem, we can leverage statistical tools that guarantee false discovery rate (FDR) requirements [24]. To elaborate, define as $\hat{\Lambda}_{\mathcal{Z}}$ the subset of hyperparameters selected by an MHT mechanism. The FDR is defined as the expected proportion of unreliable hyperparameters in set $\hat{\Lambda}_{\mathcal{Z}}$. Therefore, controlling the FDR amounts to finding a subset $\hat{\Lambda}_{\mathcal{Z}}\subseteq\Lambda$ that satisfies the inequality
where $\mathbf{1}\{\cdot\}$ is the indicator function, and the expectation is taken over the unknown data distribution $P_{Z}$. The FDR constraint (8) ensures that the average fraction of unreliable hyperparameters in set $\hat{\Lambda}_{\mathcal{Z}}$ is upper bounded by $\delta$.